# SENSEI: Semantic Exploration Guided by Foundation Models to Learn Versatile World Models

### v1

@@ 5 Discussion
We have introduced Sensei, a framework for guiding the exploration of model-based agents through foundation models. Sensei bootstraps a model of interestingsness from previously generated play data. On this dataset, Sensei prompts a VLM to compare images with respect to their interestingness and distills a semantic reward function. Sensei learns an exploration policy via model-based RL using two sources of intrinsic rewards: (1) trying to reach states with high semantic interestingness and (2) branching out from these states to maximize epistemic uncertainty. We show that in the video game environments of MiniHack and a robotic simulation, this strategy leads to more meaningful interactions, e.g. opening a chest with a key or manipulating objects on a desk. In both environments, Sensei accumulates more rewards than the state-of-the-art exploration method Plan2Explore (Sekar et al., 2020), as it solves the human-designed sparse reward tasks already during exploration.
@@
In contrast to other methods of foundation model-guided exploration (Klissarov et al., 2023; Wang et al., 2024), Sensei combines semantic exploration with world model learning, thereby learning an internal model of “interestingness.” This is a sensible design choice when working with world models (as detailed in Suppl. A.3). Previous variants of Motif (Klissarov et al., 2023; Wang et al., 2024) require the complete observation in order to compute the foundation model-based reward. After training, Sensei can predict semantic rewards from its latent states, thus, also evaluating imagined, hypothetical states. As demonstrated by our experiments, this can lead to significantly faster policy learning, since both VLM guidance as well as model-based RL improve sample efficiency.
@@
Sensei benefits from fully-observable observations, e.g. images that capture all relevant aspects of the environment. The VLM annotations, and as a result the distilled reward function, degrade when dealing with occlusions. In Robodesk, we used a second camera angle for VLM annotations to reduce noise. In future work this could be remedied further by annotating videos to better convey temporal or partially-observable information.
@@
We hypothesize that Sensei learns a versatile world model from exploration, as it is trained on a more interaction-rich dataset. In future work, we plan to further examine the versatility of the explored world model by employing it for zero-shot model-based planning  (Sancaktar et al., 2022). Furthermore, in Suppl. D.2, by comparing two sources for the initial play dataset $\mathcal{D}^{\mathrm{init}}$, we observe that Sensei partially reinforces the trends existing in the initial exploration round. To further reinforce more complex types of interactions, our reward function could be refined in a new round of VLM annotations from a Sensei run. As a result, Sensei could potentially unlock increasingly complex behavior with each generation. Besides that, photorealism of observations are likely to help VLM annotations because a large portion of VLMs’ training data comes from real world photos or videos. Thus, we think that Sensei could scale well to more realistic applications or real-world tasks.
@@ A.1 World model
Both our Plan2Explore baseline as well as our ensemble predictors (Eq. 5) are based on the re-implementation on top of DreamerV3. The most notable difference is that in original Plan2Explore the ensemble is trained to predict image encodings (Sekar et al., 2020), whereas the new version is trained to predict stochastic states $\bm{z}_{t}$. Recent re-implementations (Hafner, 2021; Hafner et al., 2022; Gumbsch et al., 2024) also used Plan2Explore with ensemble disagreement over $\bm{z}_{t}$ as a baseline and verified a strong exploration performance.
@@
In practice, we compute exploration rewards (Eq. 7) using two reward factors for each loss term
$r_{t}^{\mathrm{expl}}=\begin{cases}\alpha^{\mathrm{explore}}\hat{r}^{\mathrm{sem}}_{t}+\beta^{\mathrm{explore}}r^{\mathrm{dis}}_{t},&\text{if}\quad\hat{r}^{\mathrm{sem}}_{t}\geq Q_{k}(\hat{r}^{\mathrm{sem}});\\ \alpha^{\mathrm{go}}\hat{r}^{\mathrm{sem}}_{t}+\beta^{\mathrm{go}}r^{\mathrm{dis}}_{t},&\text{otherwise.}\end{cases}$ (8)
i.e. $\alpha$ to scale $\hat{r}^{\mathrm{sem}}_{t}$ and $\beta$ to scale $r^{\mathrm{dis}}_{t}$. When training the value function with DreamerV3, the scale of the reward sources are normalized. To compute this normalization for the exploration policy we use $\alpha^{\mathrm{explore}}$ and $\beta^{\mathrm{explore}}$ of the high percentile region of interestingness ($\geq Q_{k}$).
@@ Appendix B Hyperparameters
Robodesk KeyRoom KeyChest
@@
Rssm size M S S
Ensemble size $N$ 8 8 8
Train ratio 512 512 512
@@
Quantile $0.75$ $0.85$$0.75$ $0.80$ $0.90$ $0.90$
$\alpha^{\mathrm{explore}}$ 0.1 - 0.1 - 0.05 - 0.01 $0.3$ $0.25$
$\beta^{\mathrm{explore}}$ 1 - 1 - 1 - 1 $1$ $1$
$\alpha^{\mathrm{go}}$ 1 - 1 - 1 - 1 1 1
$\beta^{\mathrm{go}}$ 0 - 0 - 0 - 0 $0.1$ $0.05$
@@
VLM GPT-4 turbo (right) & GPT-4 omni (left) GPT-4 omni GPT-4 omni
Temperature $0.2$ $0.2$ $0.2$
Dataset size 200K 100K 100K
Image res. $224\!\times\!224$ $80\!\times\!80$ $80\!\times\!80$
@@
Batch size 32 - 64 - 32 - 32 32 32
Learning rate $10^{-5}$ $10^{-5}$- $3\!\times\!10^{-5}$ $3\!\times\!10^{-5}$ $10^{-4}$ $10^{-4}$
Weight decay $10^{-5}$ 0 - 0 - 0 $10^{-5}$ $10^{-4}$
@@
Action repeat 2 1 1
Episode length 250 600 800
Steps of exploration 1M 500K 500k
For the exploration policy in Robodesk we use different values for the four different variants tested. The values listed here stand for, from left to right: GPT-4 with Plan2Explore data using two camera angles for VLM annotations, GPT-4 with Plan2Explore data using only the right camera angle for annotations, Oracle with Plan2Explore data, and Oracle with CEE-US data (corresponding to a more interaction-rich exploration dataset for $\mathcal{D}^{\mathrm{init}}$). The VLM-Motif training hyperparameters are also listed in the same order.
@@
We run DreamerV3 with the same world model setup as Sensei and Plan2Explore.
We use an open source PPO (Schulman et al., 2017) implementation of Hafner (2024) [^4] https://github.com/danijar/embodied, version v1.2 optimized to work well across multiple environments with a fixed set of hyperparameters (details in Hafner et al., 2023, supplementary material). We build our RND (Burda et al., 2019) implementation on top of PPO. For the predictor and target network we use a ResNet with 3 convolutional layers followed by 5 dense layers. We only use the intrinsic reward to train a PPO agent. Intrinsic rewards are normalized as outlined in Burda et al. (2019). While Burda et al. (2019) also normalize input observations through a running statistics, we found that using LayerNorm at the input layer leads to slightly more interactions in Robodesk.
@@ C.1 Robodesk
Figure 8: Robodesk environment: We modify the default top-down camera view (a) to a side view with less occlusion (b).
@@
Robodesk uses only an image observation, depicting the current scene, which we scale down ($64\times 64$ pixels). However, we found that the default top-down view often had occlusions and was hard to interpret from a single image (Fig. 8(a)). Thus, we used a different camera angle showing the robot from one side (Fig. 8(b)). With this view objects and the drawer were rarely occluded; however, lights that turn on from button presses were not as visible anymore.
@@
  • Drawer: Opening the drawer fully (open_drawer) or opening it slightly (open_drawer_light). We introduced the latter task.
@@ C.2 MiniHack
C.3 VLM prompting
@@ C.3.1 Robodesk
Figure 10: Left camera observation in Robodesk that is used only during annotation with GPT-4.
Due to occlusions, we annotate the same pair from the initial dataset $\mathcal{D}^{\mathrm{init}}$ with the same prompt using images from two camera angles: right (Fig. 8(b)) and left (Fig. 10). A pair is deemed valid only if the GPT-4 response is the same across both camera angles, otherwise the pair and the annotation are removed from the dataset. The VLM-Motif training as well as the world model training are still executed using only the right camera image, such that during free play with Sensei we only rely on the right camera images as input.
@@
C.3.2 Robodesk: Zero Pre-Knowledge General Prompt
To demonstrate the generality of Sensei, we also test a version with a general prompt (Sensei general), using a multi-turn strategy for getting the annotations for a pair of images. This version of the prompt does not contain any external environment specific knowledge. First, we show a picture from the robotic environment (shown in Fig. 8(b)) and ask the VLM (gpt-4o-2024-05-13) to describe what it sees in the scene using the following prompt:
@@ C.3.3 MiniHack
C.4 Oracle for Interestingness
In Robodesk, we also use an Oracle of interestingness to annotate the pairs as an ablation (see Suppl. D.2). Our goal here is to showcase an upper-bound of performance on Sensei without the noisiness of VLMs. For the Oracle, we deem a state interesting if: (1) any one of the entities are in motion (here only for the ball we make an exception that the ball should be in motion with the end effector close to it as the ball in the environment is unimpeded by friction), (2) if the drawer is opened, (3) if the drawer/sliding cabinet is not yet in motion, but the end effector is very close to their handles, (4) if the upright and flat blocks are not yet in motion but the end effector is very close to them (almost touching), (5) if the stacking task is solved. With these statements, we essentially cover the range of tasks defined in the Robodesk environment, as they are shown in Fig. 14.
Appendix D Extended Results
D.1 Minihack: Extended Results
@@
D.2 Robodesk: Sensei Ablations
In Robodesk, we compare different versions of Sensei in order to analyze the effect of the VLM and the initial exploration data on Sensei performance (Fig. 12). First, we showcase Sensei results when annotating the initial exploration dataset from Plan2Explore with only the right camera images. In this case, we use the whole 200K pairs in the dataset, without any pruning. In another ablation, we replace the VLM (GPT-4) with a hand-crafted Oracle (see Suppl. C.4 for how the oracle is computed) for annotating the pairs. After the oracle annotations, we distill these preferences into VLM-Motif for Sensei, following the same procedure as before. Furthermore, we compare two initial datasets $\mathcal{D}^{\mathrm{init}}$ of self-supervised exploration collected either by CEE-US (Sancaktar et al., 2022) or by Plan2Explore for the oracle Sensei versions. CEE-US uses vector-based position of entities for information-gain-based exploration, in comparison to Plan2Explore, which works on the pixel-level. Due to the privileged inputs, $\mathcal{D}^{\mathrm{init}}_{\mathrm{CEE-US}}$ contains more complex interactions. We compare 1M steps of exploration with the four versions of Sensei and Plan2Explore.

### v2

@@ 5 Discussion
We have introduced Sensei, a framework for guiding the exploration of model-based agents through foundation models. Sensei bootstraps a model of interestingness from previously generated play data. On this dataset, Sensei prompts a VLM to compare images with respect to their interestingness and distills a semantic reward function. Sensei learns an exploration policy via model-based RL using two sources of intrinsic rewards: (1) trying to reach states with high semantic interestingness and (2) branching out from these states to maximize epistemic uncertainty. We show that in the video game environments of MiniHack and Pokémon Red and a robotic simulation, this strategy leads to more meaningful interactions, e.g. opening a chest with a key or manipulating objects on a desk.
@@
Unlike prior work of foundation model-guided exploration (Klissarov et al., 2023; Wang et al., 2024), Sensei learns an internal model of interestingness. This is a sensible design choice when working with world models (as detailed in Suppl. A.3), enabling Sensei to predict semantic rewards also while imagining states during policy training. We demonstrate that this can lead to significantly faster learning, since both VLM guidance as well as model-based RL improve sample efficiency.
@@
Sensei benefits from fully-observable observations, e.g. images that capture all relevant aspects of the environment. The VLM annotations degrade when dealing with occlusions. In Robodesk we mitigate this using multiple camera angles. In future work this could be remedied further by annotating videos to better convey temporal or partially-observable information.
@@
One promising direction is to systematically investigate iterative refinement of the semantic reward function. In Suppl. E.9, we show that incorporating Sensei-collected data into a second round of annotations reduces out-of-distribution errors for states not present in the initial annotation set. We believe this iterative process can unlock increasingly complex behaviors with each generation. Another avenue is to explore Sensei in photorealistic or real-world environments. Photorealism of observations are likely to help VLM annotations because a large portion of VLMs’ training data comes from real world photos or videos. Thus, Sensei is likely to scale well to these settings.
Acknowledgements
@@ Appendix A SENSEI: Implementation Details
We provide code with VLM-Motif checkpoints at https://github.com/martius-lab/sensei.
@@ A.1 World model
Both our Plan2Explore baseline as well as our ensemble predictors (Eq. 5) are based on the re-implementation on top of DreamerV3. The most notable difference is that in original Plan2Explore the ensemble is trained to predict image encodings (Sekar et al., 2020), whereas the new version is trained to predict stochastic states $\bm{z}_{t}$. Recent re-implementations (Hafner, 2021; Hafner et al., 2022; Gumbsch et al., 2024) also used Plan2Explore with ensemble disagreement over $\bm{z}_{t}$ as a baseline and verified a strong exploration performance. For our experiments on task-based exploration (Sec. 4.4), we use a weighted sum of extrinsic and intrinsic rewards, i.e. $r_{t}+\alpha r^{\mathrm{dis}}_{t}$. We determined the best value $\alpha=0.5$ through a hyperparameter search, as for Sensei’s hyperparameters.
@@
In practice, we compute exploration rewards (Eq. 7) overall five reward factors:
$r_{t}^{\mathrm{expl}}=\chi r_{t}+\begin{cases}\alpha^{\mathrm{explore}}\hat{r}^{\mathrm{sem}}_{t}+\beta^{\mathrm{explore}}r^{\mathrm{dis}}_{t},&\text{if}\quad\hat{r}^{\mathrm{sem}}_{t}\geq Q_{k}(\hat{r}^{\mathrm{sem}});\\ \alpha^{\mathrm{go}}\hat{r}^{\mathrm{sem}}_{t}+\beta^{\mathrm{go}}r^{\mathrm{dis}}_{t},&\text{otherwise.}\end{cases}$ (8)
i.e. $\chi\in\{0,1\}$ to scale extrinsic rewards $r_{t}$, $\alpha$ to scale semantic rewards $\hat{r}^{\mathrm{sem}}_{t}$ and $\beta$ to scale uncertainty-based rewards $r^{\mathrm{dis}}_{t}$. We set $\chi=0$ for task-free exploration and $\chi=1$ for task-based exploration (Sec. 4.4). When training the value function with DreamerV3, the scale of the reward sources are normalized. To compute this normalization for the exploration policy we use $\alpha^{\mathrm{explore}}$ and $\beta^{\mathrm{explore}}$ of the high percentile region of interestingness ($\geq Q_{k}$).
@@ Appendix B Hyperparameters
Robodesk KeyRoom KeyChest Pokémon Red
@@
Rssm size M S S L
Ensemble size $N$ 8 8 8 8
Train ratio 512 512 512 512
@@
Quantile $0.75$ $0.85$$0.75$ $0.80$ $0.90$ $0.90$ $0.6$
$\chi$ 0 - 0 - 0 - 0 00 00 $1$
$\alpha^{\mathrm{explore}}$ 0.1 - 0.1 - 0.05 - 0.01 $0.3$ $0.25$ $0.025$
$\beta^{\mathrm{explore}}$ 1 - 1 - 1 - 1 $1$ $1$ $0.5$
$\alpha^{\mathrm{go}}$ 1 - 1 - 1 - 1 1 1 $0.5$
$\beta^{\mathrm{go}}$ 0 - 0 - 0 - 0 $0.1$ $0.05$ $0.1$
@@
VLM GPT-4 turbo (right) & GPT-4 omni (left) GPT-4 omni GPT-4 omni GPT-4 omni
Temperature $0.2$ $0.2$ $0.2$ $0.2$
Dataset size 200K 100K 100K 100K
Image res. $224\!\times\!224$ $80\!\times\!80$ $80\!\times\!80$ $1202\!\times\!1080$
@@
Batch size 32 - 64 - 32 - 32 32 32 32
Learning rate $10^{-5}$ $10^{-5}$- $3\!\times\!10^{-5}$ $3\!\times\!10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-5}$
Weight decay $10^{-5}$ 0 - 0 - 0 $10^{-5}$ $10^{-4}$ 00
@@
Action repeat 2 1 1 1
Episode length 250 600 800 1k/2k/4k
Steps of exploration 1M 500K 500k 750k
For the exploration policy in Robodesk we use different values for the four different variants tested. The values listed here stand for, from left to right: GPT-4 with Plan2Explore (P2X) data using two camera angles for VLM annotations, GPT-4 with P2X data using only the right camera angle, Oracle with P2X data, and Oracle with CEE-US data (corresponding to a more interaction-rich exploration dataset $\mathcal{D}^{\mathrm{init}}$). The VLM-Motif hyperparameters are also listed in the same order.
@@
We run DreamerV3 with the same world model setup as Sensei and Plan2Explore. We use an open source PPO (Schulman et al., 2017) implementation of Hafner (2024) [^4] https://github.com/danijar/embodied, version v1.2 optimized to work well across multiple environments with a fixed set of hyperparameters (details in Hafner et al., 2023, supplementary material). We build our RND (Burda et al., 2019) implementation on top of PPO. For the predictor and target network we use a ResNet with 3 convolutional layers followed by 5 dense layers. We only use the intrinsic reward to train a PPO agent. Intrinsic rewards are normalized as outlined in Burda et al. (2019). While Burda et al. (2019) also normalize input observations through a running statistics, we found that using LayerNorm at the input layer leads to slightly more interactions in Robodesk.
@@ C.1 Robodesk
(c) Left camera
Figure 9: Robodesk environment: We modify the default top-down camera view (a) to a side view with less occlusion (b). For annotation with GPT-4 we also provide a left camera observation (c).
@@
Robodesk uses only an image observation, depicting the current scene, which we scale down ($64\times 64$ pixels). However, we found that the default top-down view often had occlusions and was hard to interpret from a single image (Fig. 9(a)). Thus, we used a different camera angle showing the robot from one side (Fig. 9(c)). With this view objects and the drawer were rarely occluded; however, lights that turn on from button presses were not as visible anymore.
@@
  • Drawer: Opening the drawer fully (open_drawer), opening the drawer half-way (open_drawer_medium), or opening it slightly (open_drawer_light). We introduced the latter tasks.
@@ C.2 MiniHack
C.3 Pokémon Red
(a) map navigation
(b) battle screen
(c) low-resolution observation
Figure 11: Pokémon Red poses a strong exploration challenge as it requires an agent to learn (1) to navigate a complex overworld map (a) and (2) to battle and catch Pokémon (b). We only use the down-scaled game screen image as observations (c).
We evaluate SENSEI’s ability for semantic exploration in the classic Game Boy game, Pokémon Red. Pokémon Red presents an extremely challenging exploration problem due to: (1) its vast and interconnected world, composed of towns, routes, forests, and other areas that must be navigated, and (2) its complex battle system, which requires semantic knowledge to understand type interactions (e.g., Water attacks are strong against Fire-type Pokémon). Thus, an agent needs strong exploration capabilities to progress in the game’s primary objective, i.e., becoming the Pokémon Champion by defeating Gym Leaders and collecting their badges. We base our implementation on PokeGym v0.1.1 (Suarez, 2024), which is based on Whidden (2023), with minor modifications as detailed below.
Observations
Unlike previous RL agents applied to Pokémon (Whidden, 2023; Pleines et al., 2025), we provide only the raw game screen image as input, without access to the internal game state or additional memory. When input into the world model, we downscale the game screen (to 64x64 pixels) in order to save compute (see Fig. 11(c)). For VLM annotations, we use the original size (1202x1080 pixels), such that all text is clearly readable (see Fig. 11(b)).
Actions
The agent controls the game using a 6-dimensional action space corresponding to Game Boy button presses (Left, Right, Up, Down, A, B). Since the game only advances upon button presses (except during attack animations) we only apply an action every 1.5 seconds real time game play (frame skip of 96) to manage episode length.
Rewards
The agent receives a weighted sum of rewards $r_{t}=\sum\phi^{i}r_{t}^{i}$ for different in-game events $i$. Rewarded events include leveling up, catching Pokémon, encountering strong opponents, healing Pokémon, visiting new map tiles, earning badges, and a penalty for blacking out after losing a battle. We leave the default values $\phi^{i}$ from Suarez (2024) except we increase $\phi^{\mathrm{tiles}}=0.1$ for reaching new map tiles (previously set to $0.01$), as we found this improves exploration.
Episode Length
To scaffold exploration, we gradually increase the maximum episode length over environment steps: 1k length for 0–25k steps, 2k length for 25–50k, 4k length for 50–75k, and 8k length when continuing exploration beyong 75k steps.
Starting point
As in Whidden (2023), we skip the initial “tutorial” phase of the game, and start only when the player can freely move and catch Pokémon. In-game, this corresponds to the point after delivering Oak’s Parcel and receiving the Pokédex and Poké Balls from Professor Oak. We use the same checkpoint as Whidden (2023), starting with a level 6 Squirtle named AAAAAAAAAA.
Map segments
The accessible game world in Pokémon Red prior to defeating the first Gym is already expansive, comprising a variety of interconnected areas. Many of these are distinct maps that load separately when the player enters or exits a building (see Fig. 12). To evaluate spatial exploration, we segment this world into discrete map segments, as illustrated in Fig. 8(a). Each route and town is treated as a separate segment, as are buildings or enclosed areas like forests that load a dedicated sub-map. This amounts to 25 map segments that can be accessed before beating the first Pokémon Gym. Reaching the first Pokémon Gym requires navigating through 10 such segments, which we enumerate in order of appearance (cf. Fig. 8(a)):
Number Name Type
0 Oak’s Lab building
1 Pallet Town town
2 Route 1 route
3 Viridian City town
4 Route 2 route
5 Viridian Forest South Gate building
6 Viridian Forest forest
7 Viridian Forest North Gate building
8 Pewter City town
9 Pewter Gym building
This segmentation allows us to quantify exploration progress by measuring the highest-indexed map segment reached during an episode.
Figure 12: Full map of Pokémon Red accessible prior to defeating the first Gym. If entering or exiting a building brings the agent to a submap, this is indicated by a line. The map is modified from https://blog.vjeux.com/2023/project/pokemon-red-blue-map.html
Appendix D VLM prompting
@@ D.1 Robodesk
Due to occlusions, we annotate the same pair from the initial dataset $\mathcal{D}^{\mathrm{init}}$ with the same prompt using images from two camera angles: right (Fig. 9(c)) and left (Fig. 9(c)). A pair is deemed valid only if the GPT-4 response is the same across both camera angles, otherwise the pair and the annotation are removed from the dataset. The VLM-Motif training as well as the world model training are still executed using only the right camera image, such that during free play with Sensei we only rely on the right camera images as input.
@@
D.2 Robodesk: Zero Pre-Knowledge General Prompt
To demonstrate the generality of Sensei, we also test a version with a general prompt (Sensei general), using a multi-turn strategy for getting the annotations for a pair of images. This version of the prompt does not contain any external environment specific knowledge. First, we show a picture from the robotic environment (shown in Fig. 9(c)) and ask the VLM (gpt-4o-2024-05-13) to describe what it sees in the scene using the following prompt:
@@ D.3 MiniHack
D.4 Pokémon Red
For Pokémon Red we set the goal to reach the first Pokémon Gym (the first boss battle of the game): We assume that GPT-4 (gpt-4o-2024-05-13) was extensively trained on game play data and various walkthroughs of Pokémon Red and contains sufficient knowledge of the game. Thus, we again use a multi-turn strategy for image annotations, as with Sensei general (see Suppl. D.2), where we first ask the VLM for a game play description, which we then use as context for further annotations. We first asked the VLM:
Your task is to help me play the Game Boy game Pokémon Red. I just obtained my starter Pokémon, Squirtle. My goal is to find and defeat the Gym Leader, Brock. What do I need to do, and which areas do I need to traverse to achieve this goal? Keep it very short.
We generated five different responses from GPT-4. We sample from them uniformly as context for image-based comparisons. We provide two observation images with the following prompt:
Here are two screenshots from the game. Which image depicts a game state that is closer to my goal? Answer in maximum one word: 0 if the first state is better, 1 if the second state is better.
D.5 Oracle for Interestingness
In Robodesk, we also use an Oracle of interestingness to annotate the pairs as an ablation (see Suppl. E.2). Our goal here is to showcase an upper-bound of performance on Sensei without the noisiness of VLMs. For the Oracle, we deem a state interesting if: (1) any one of the entities are in motion (here only for the ball we make an exception that the ball should be in motion with the end effector close to it as the ball in the environment is unimpeded by friction), (2) if the drawer is opened, (3) if the drawer/sliding cabinet is not yet in motion, but the end effector is very close to their handles, (4) if the upright and flat blocks are not yet in motion but the end effector is very close to them (almost touching), (5) if the stacking task is solved. With these statements, we essentially cover the range of tasks defined in the Robodesk environment, as they are shown in Fig. 16.
Appendix E Extended Results
E.1 MiniHack: Extended Results
@@
E.2 Robodesk: Sensei Ablations
In Robodesk, we compare different versions of Sensei in order to analyze the effect of the VLM and the initial exploration data on Sensei performance (Fig. 14). First, we showcase Sensei results when annotating the initial exploration dataset from Plan2Explore with only the right camera images. In this case, we use the whole 200K pairs in the dataset, without any pruning. In another ablation, we replace the VLM (GPT-4) with a hand-crafted Oracle (see Suppl. D.5 for how the oracle is computed) for annotating the pairs. After the oracle annotations, we distill these preferences into VLM-Motif for Sensei, following the same procedure as before. Furthermore, we compare two initial datasets $\mathcal{D}^{\mathrm{init}}$ of self-supervised exploration collected either by CEE-US (Sancaktar et al., 2022) or by Plan2Explore for the oracle Sensei versions. CEE-US uses vector-based position of entities for information-gain-based exploration, in comparison to Plan2Explore, which works on the pixel-level. Due to the privileged inputs, $\mathcal{D}^{\mathrm{init}}_{\mathrm{CEE-US}}$ contains more complex interactions. We compare 1M steps of exploration with the four versions of Sensei and Plan2Explore.